# Beyond the Surface: Probing the Ideological Depth of Large Language Models

### v1

@@ Appendix A Factor Analysis of LLM Responses
Figure 15: Scree Plot of Factor Eigenvalues
Factor Eigenvalue Proportion Cumulative
@@
Figure 16: Eigenvalue Table (Unrotated Setting)
@@
Factor Variance Proportion Cumulative
@@
Table 3: Rotated Factor Variance Table
@@
An examination of the factor loadings confirms that this primary dimension aligns conceptually with the traditional left/right economic spectrum, with items related to [e.g., ‘Political & Ideological Stances’, ‘Tax Policy’] loading strongly onto it. The subsequent factors appear to capture more nuanced, secondary themes. Therefore, while acknowledging the existence of other dimension, we justify focusing our evaluation on this primary dimension.
Appendix B Validating the Multidimensional Item Response Theory Model
A central challenge in Multidimensional Item Response Theory (MIRT) models is the issue of identifiability. The latent ideological space is invariant to rotation, reflection, and translation, meaning that an infinite number of parameter sets ($\boldsymbol{\theta},\boldsymbol{\alpha},\beta$) can produce the identical likelihood. Without proper constraints, this leads to unstable parameter estimates and prevents meaningful interpretation of the dimensions. This appendix documents our iterative approach to resolving this issue. The success of each identification strategy is evaluated by its ability to produce stable and externally valid ideal points, which we assess by correlating our estimates with the widely-used DW-NOMINATE [21] and IDEAL scores [8].
B.1 Model Specification
@@
B.2 Evolution of Identification Strategy
@@
Figure 17: Model Performance by Identification Strategy.
B.2.1 Identification via Priors (No Fixed References)
@@
B.2.2 Two-Point Constraint Identification
Following the precedent set by the IDEAL model by Clinton et al. [8], we introduced constraints by fixing the positions of reference candidates with well-established, opposing ideologies to anchor the primary dimension. For instance, a strongly liberal candidate was fixed at $\theta_{j_{1},1}=-2.0$ and a strongly conservative candidate was fixed at $\theta_{j_{2},1}=2.0$. This successfully stabilized the first dimension, yielding consistently high correlations with DW-NOMINATE and IDEAL. However, the second dimension remained unidentified. While the plane of candidate ideal points was fixed relative to the two reference points, it could still rotate around the axis connecting them, leading to highly variable and unreliable estimates for the second dimension across repeated trials. This can be seen in the middle of Figure 17 where Dim. 1 becomes highly correlated and very stable (small bars), but Dim. 2 remains uncertain.
B.2.3 Three-Point Constraint Identification
@@
Appendix C Instruction Prompts For Adopting Different Personas
C.1 Liberal Persona
@@ C.2 Conservative Persona
Appendix D Promprs Used In Feature Qualitative Analysis
D.1 Predictive Validity

### v2

@@ Appendix E Factor Analysis of LLM Responses
Figure 13: Scree plot of eigenvalues from factor analysis of model responses, showing one dominant ideological dimension.
Factor Eigenvalue Prop. Cumulative
@@
Table 4: Unrotated factor eigenvalues. The first factor explains most of the variance, justifying a single-axis interpretation.
@@
Factor Variance Prop. Cumulative
@@
Table 5: Rotated factor variances after Varimax rotation. The first rotated factor remains dominant, capturing the main ideological dimension.
@@
An examination of the factor loadings confirms that this primary dimension aligns conceptually with the traditional left/right economic spectrum, with items related to [e.g., ‘Political & Ideological Stances’, ‘Tax Policy’] loading strongly onto it. The subsequent factors appear to capture more nuanced, secondary themes. Therefore, while acknowledging the existence of other dimensions, we justify focusing our evaluation on this primary dimension.
Appendix F Validating the Multidimensional Item Response Theory Model
A central challenge in Multidimensional Item Response Theory (MIRT) models is the issue of identifiability. The latent ideological space is invariant to rotation, reflection, and translation, meaning that an infinite number of parameter sets ($\boldsymbol{\theta},\boldsymbol{\alpha},\beta$) can produce the identical likelihood. Without proper constraints, this leads to unstable parameter estimates and prevents meaningful interpretation of the dimensions. This appendix documents our iterative approach to resolving this issue. The success of each identification strategy is evaluated by its ability to produce stable and externally valid ideal points, which we assess by correlating our estimates with the widely-used DW-NOMINATE poole1985spatial and IDEAL scores clinton2004statistical.
F.1 Model Specification
@@
F.2 Evolution of Identification Strategy
@@
Figure 14: Correlation of identification strategies in multidimensional IRT. Three-point anchoring yields the most stable and interpretable ideological space.
F.2.1 Identification via Priors (No Fixed References)
@@
F.2.2 Two-Point Constraint Identification
Following the precedent set by the IDEAL model by clinton2004statistical, we introduced constraints by fixing the positions of reference candidates with well-established, opposing ideologies to anchor the primary dimension. For instance, a strongly liberal candidate was fixed at $\theta_{j_{1},1}=-2.0$ and a strongly conservative candidate was fixed at $\theta_{j_{2},1}=2.0$. This successfully stabilized the first dimension, yielding consistently high correlations with DW-NOMINATE and IDEAL. However, the second dimension remained unidentified. While the plane of candidate ideal points was fixed relative to the two reference points, it could still rotate around the axis connecting them, leading to highly variable and unreliable estimates for the second dimension across repeated trials. This can be seen in the middle of Figure 14 where Dim. 1 becomes highly correlated and very stable (small bars), but Dim. 2 remains uncertain.
F.2.3 Three-Point Constraint Identification
@@
Appendix G Instruction Prompts For Adopting Different Personas
G.1 Liberal Persona
@@ G.2 Conservative Persona
Appendix H Promprs Used In Feature Qualitative Analysis
H.1 Predictive Validity